# Automata Learning from Preference and Equivalence Queries

### v2

@@ Abstract
Active automata learning from membership and equivalence queries is a foundational problem with numerous applications. We propose a novel variant of the active automata learning problem: actively learn finite automata using preference queries—i.e., queries about the relative position of two sequences in a total order—instead of membership queries. Our solution is Remap, a novel algorithm which leverages a symbolic observation table along with unification and constraint solving to navigate a space of symbolic hypotheses (each representing a set of automata), and uses satisfiability-solving to construct a concrete automaton (specifically a Moore machine) from a symbolic hypothesis. Remap is guaranteed to correctly infer the minimal automaton with polynomial query complexity under exact equivalence queries, and achieves PAC–identification ($\varepsilon$-approximate, with high probability) of the minimal automaton using sampling-based equivalence queries. Our empirical evaluations of Remap on the task of learning reward machines for two reinforcement learning domains indicate Remap scales to large automata and is effective at learning correct automata from consistent teachers, under both exact and sampling-based equivalence queries.
Introduction
@@
Active automata learning has applications from software engineering (Schuts, Hooman, and Vaandrager 2016; Aarts et al. 2012) and verification (Lin et al. 2014) to interpretable machine learning (Weiss, Goldberg, and Yahav 2019) and learning reward machines (Tappler et al. 2019; Gaon and Brafman 2020; Xu et al. 2021; Dohmen et al. 2022). The classical problem formulation involves a teacher with access to a regular language and a learner which asks membership and equivalence queries to infer a finite automaton describing the regular language (Angluin 1987; Gold 1978).
Consider an alternative formulation: learning a finite automaton from preference and equivalence queries [^1] Shah et al. (2023) primarily investigates choosing between membership and preference queries. A preference query resolves the relative position of two sequences in a total order available to the teacher. Learning from preference data has applications in fine-tuning language models (Ouyang et al. 2022), learning conditional preference networks (Koriche and Zanuttini 2009; Guerin, Allen, and Goldsmith 2013), inferring reinforcement learning policies (Christiano et al. 2017), and learning Markovian reward functions (Bıyık, Talati, and Sadigh 2022; Sadigh et al. 2017; Bewley and Lécué 2021; Kalra and Brown 2023).
@@
Applications for learning finite automata from preferences over sequences include inferring sequence classifications (e.g. program executions (Giannakopoulou, Rakamaric, and Raman 2012), vehicle maneuvers, human-robot interactions) using ordered classes (e.g. safe, risky, dangerous, fatal), learning preference models (Herin, Perny, and Sokolovska 2023, 2024), distilling interpretable preference models (Weiss, Goldberg, and Yahav 2019), and inferring reward machines. However, no method currently exists for learning finite automata from preferences with the termination and correctness guarantees enjoyed by classical automata learning algorithms such as $\textrm{L}^{*}$ (Angluin 1987).
@@
Remap (Figure 1) features termination and correctness guarantees for exact and probably approximately correct (Valiant 1984) (PAC) identification (Angluin 1988) of the desired automaton, with a strong learner capable of symbolic reasoning and constraint-solving to offset the weaker preference-based signals from the teacher. By using unification to navigate the symbolic space of hypotheses, and constraint-solving to construct a concrete automaton from a symbolic hypothesis, Remap identifies, in a polynomial number of queries, the minimal Moore machine isomorphic to one describing the teacher’s total order when using exact equivalence queries. Using sampling-based equivalence queries achieves PAC–identification. Our empirical evaluations apply Remap to learn reward machines for sequential-decision making domains in the reward machine literature. We measure query complexity for the exact and PAC–identification settings, and measure empirical correctness for the PAC–identification setting.
@@
We contribute (a) Remap, a novel $\textrm{L}^{*}$ style algorithm for learning Moore machines using preference and equivalence queries; under exact and PAC-identification settings we provide (b) theoretical analysis of query complexity, correctness, and minimalism, and (c) supporting empirical results, demonstrating the efficacy of the algorithm’s ability to learn reward machines from preference and equivalence queries.
Background
Figure 2: Differences between Remap and $\textrm{L}^{*}$. Full example: Appendix, Figure 6. In (a) $\textrm{L}^{*}$ records concrete values from membership queries in its observation table, whereas (c,d,f) Remap uses SymbolicFill to perform preference queries and record variables in its symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$. Part (c) initializes $\langle S,E,T;\mathcal{C},\Gamma\rangle$; (d) expands $S$ with sequence $b$, followed by SymbolicFill, yielding a unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$; (e) MakeHypothesis yields concrete hypothesis $h_{1}$ from the symbolic hypothesis and constraints in $\mathcal{C}$; (f) submits $h_{1}$ via equivalence query, receives and processes counterexample $bab$ with feedback $f(bab)=0$ from teacher by adding $bab$, $ba$, and $b$ to $S$, performs a SymbolicFill, sets the value of equivalence class for $T(bab)$ to $f(bab)$, yielding an inconsistent table. Part (b) shows the ground truth Moore machine with $\Sigma^{I}=\{a,b\},\Sigma^{O}=\{0,1\}$. Green symbolizes new information. Other colors in (f) visually highlight inconsistencies.
@@
Prior to introducing Remap, we provide preliminaries of finite automata, followed by a discussion of Angluin’s $\textrm{L}^{*}$ algorithm in order to highlight how Remap differs from $\textrm{L}^{*}$.
@@
Finite Automata  Automata describe sets of sequences. An alphabet $\Sigma$ is a set whose elements can be used to construct sequences; $\Sigma^{*}$ represents the set of sequences of any length created from elements of $\Sigma$. A sequence $s\in\Sigma^{*}$ has integer length $|s|\geq 0$; if $|s|=0$, then $s$ is the empty sequence $\varepsilon$. An element $\sigma\in\Sigma$ has length 1; if $s,t\in\Sigma^{*}$, then $s\cdot t$ represents $s$ concatenated with $t$, with length $|s|+|t|$. Different types of finite automata have different semantics. Deterministic automata feature a deterministic transition function $\delta$ defined over a set of states $Q$ and an input alphabet $\Sigma^{I}$. An output alphabet $\Sigma^{O}$ may be present to label states or transitions using a labeling function $L$. The tuple $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ is a Moore machine (Moore 1956), where $q_{0}\in Q$ is the initial state, $\delta:Q\times\Sigma^{I}\rightarrow Q$ describes transitions, and $L:Q\rightarrow\Sigma^{O}$ associates outputs with states. Extended, $\delta:Q\times(\Sigma^{I})^{*}\rightarrow Q$, where $\delta(q,\varepsilon)=q$ and $\delta(q,\sigma\cdot s)=\delta(\delta(q,\sigma),s)$. In Mealy machines (Mealy 1955), outputs are associated with transitions instead, so $L:Q\times\Sigma^{I}\rightarrow\Sigma^{O}$. Mealy and Moore machines are equivalent (Fleischner 1977) and can be converted between one another. Reward machines are an application of Mealy machines in reinforcement learning and are used to express a class of non-Markovian reward functions.
@@
In Angluin’s seminal active learning algorithm, $\textrm{L}^{*}$ (Angluin 1987), the learner learns $\hat{f}$ as a binary classifier, where $|\Sigma^{O}|=2$, to determine sequence membership of a regular language by querying $\mathcal{T}$ with: (i) membership queries, where the learner asks $\mathcal{T}$ for the value of $f(s)$ for a particular sequence $s$, and (ii) equivalence queries, where the learner asks $\mathcal{T}$ to evaluate whether for all $s\in(\Sigma^{I})^{*}$, $\hat{f}(s)=f(s)$. For the latter query, $\mathcal{T}$ returns True if the statement holds; otherwise a counterexample $c$ for which $\hat{f}(c)\neq f(c)$ is returned. An observation table $\langle S,E,T\rangle$ records the concrete observations acquired by the learner’s queries (see Figure 2a). Here, $S$ is a set of prefixes, $E$ is a set of suffixes, and $T$ is the empirical observation function that maps sequences to output values—$T:(S\cup(S\cdot\Sigma^{I}))\cdot E\rightarrow\Sigma^{O}$. The observation table is a two-dimensional array; $s\in(S\cup(S\cdot\Sigma^{I}))$ indexes rows, and $e\in E$ indexes columns, with entries given by $T(s\cdot e)$. Any proposed hypothesis $\hat{f}$ must be consistent with $T$. The algorithm operates by construction: a deterministic transition function must be found exhibiting consistency (deterministic transitions) and operates in a closed manner over the set of states. If the consistency or closure requirements are violated, then membership queries are executed to expand the observation table. Once a suitable transition function is found, a hypothesis $\hat{f}$ can be made and checked via the equivalence query. The algorithm terminates if $\hat{f}(s)=f(s)$ for all $s\in(\Sigma^{I})^{*}$; otherwise $\textrm{L}^{*}$ continues on by adding counterexample $c$ and all its prefixes to the table, and finds another transition function satisfying the consistency and closure requirements.
@@
Consequently, we consider how $\textrm{L}^{*}$-style learning can be used to learn Moore machines from preference queries over sequences, as a foray into understanding how the structure of finite automata can be learned from comparison information.
Problem Statement
@@
We consider how a Moore machine $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ can be actively learned from preference queries over $(\Sigma^{I})^{*}$. We focus on the case of a finite sized $\Sigma^{I}$ and $\Sigma^{O}$ known to both teacher and learner. With a preference query, the learner asks the teacher $\mathcal{T}$ which of two sequences $s_{1}$ and $s_{2}$ is preferred, or if both are equally preferable. This requires a preference model, which we assume represents a total ordering over $(\Sigma^{I})^{*}$; we also assume $\Sigma^{O}$ is totally ordered. Thus, we consider a preference model for $\mathcal{T}$ where $f:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$ is consistent with both orderings, i.e., $\mathcal{T}$ prefers $s_{1}$ over $s_{2}$ if $f(s_{1})>f(s_{2})$, or otherwise has equal preference if $f(s_{1})=f(s_{2})$.
@@
Under this problem formulation, several options for evaluating hypothesis equivalence (is $\hat{f}\equiv f$?) can be defined. The standard option requires exact equivalence: $\hat{f}(s)=f(s)$ for all $s$ in $(\Sigma^{I})^{*}$. An alternative requires the total ordering implied by $\hat{f}$ and $f$ to match. That is, for all $s,t$ in $(\Sigma^{I})^{*}$, both $f(s)\mathbf{R}_{s,t}f(t)$ and $\hat{f}(s)\mathbf{R}_{s,t}\hat{f}(t)$ are satisfied, where $\mathbf{R}_{s,t}\in\{=,>,<\}$. Since a hypothesis $\hat{f}$ satisfying the standard definition always satisfies the alternative, we focus on exact equivalence queries where $\mathcal{T}$ returns feedback along with counterexample $c$. Feedback strength directly impacts how many hypotheses per counterexample can be eliminated by the learner: returning $\hat{f}(c)\neq f(c)$ is weak feedback compared to returning $\hat{f}(c)=f(c)$ or returning $\hat{f}(c)$ is in a subset of $\Sigma^{O}$. Although Remap outputs a concrete Moore machine, Remap navigates symbolic Moore machine222i.e., a Moore machine with symbolic values as outputs. space using solely preference information, while concrete information assists in selecting the concrete hypothesis. Thus, our theoretical analysis considers equivalence queries that provide counterexamples with strong feedback.
The Remap Algorithm
@@
Central to Remap is a new construct called a symbolic observation table, shown in Figure 2c, 2d, and 2f, as well as an associated algorithm for unification (Robinson 1965) inspired by Martelli and Montanari (1976) to contain the fresh variable explosion; both enable the learner to generate symbolic hypotheses purely from observed symbolic constraints, along with a constraint solver for obtaining a concrete hypothesis. We first discuss these core components of Remap and how they fit together, then illustrate the correctness and termination guarantees.
Symbolic Observation Tables
@@
Recall that in classic $\textrm{L}^{*}$, the observation table entries are concrete values obtained from membership queries (Figure 2a). However, observations obtained from preference queries are constraints, rather than concrete values, indicating that for a pair of sequences $s_{1}$ and $s_{2}$, one of $f(s_{1})>f(s_{2})$, $f(s_{1})<f(s_{2})$, or $f(s_{1})=f(s_{2})$ holds. We therefore introduce the symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$, where $S$ is a set of prefixes, $E$ is a set of suffixes, $\mathcal{C}$ is the set of known constraints, $\Gamma$ is a context which uniquely maps sequences to variables, with the set of variables $\mathcal{V}$ in the context $\Gamma$ given by $\mathcal{V}=\{\Gamma[s\cdot e]|s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E\}$, and $T:(S\cup(S\cdot\Sigma^{I}))\cdot E\rightarrow\Gamma$ maps queried sequences to variables. Thus, in a symbolic observation table, the entry for each prefix-suffix pair is a variable, rather than a concrete value; and the constraints over those variables are stored in $\mathcal{C}$ (Figure 2c, 2d, 2f). The constraints from the preferences of $\mathcal{T}$ over $f$ about $s_{1}$ and $s_{2}$ correspond with $T(s_{1})>T(s_{2})$, $T(s_{1})<T(s_{2})$, or $T(s_{1})=T(s_{2})$, respectively.
Definition 1.
@@
An equivalence class $\mathbb{C}$ of variables is a set with the property that all members of $\mathbb{C}$ are equivalent to each other. The representative .$\mathbb{C}.\text{{rep}}$ of an equivalence class $\mathbb{C}$ is a deterministically elected member of $\mathbb{C}$. The set of variables $\mathcal{V}$ can be partitioned into disjoint equivalence classes. The set of equivalence classes $\mathcal{C}_{EC}$ corresponds with the partitioning of $\mathcal{V}$ into the smallest possible number of equivalence classes consistent with equality constraints. Let ={.$\mathcal{R}=\{\mathbb{C}.\text{{rep}}|\mathbb{C}\in\mathcal{C}_{EC}\}$ be the set of representatives.
Unification and Constraints
@@
Since preference queries return observations comparing the values of two sequences, we leverage a simple unification algorithm to ensure the number of unique variables in the table remains bounded by $|\Sigma^{O}|$. Whenever we observe the constraint $f(s_{1})=f(s_{2})$, this implies $T(s_{1})=T(s_{2})$, so we add $\Gamma[s_{1}]$ and $\Gamma[s_{2}]$ to an equivalence class $\mathbb{C}$, and elect a representative from $\mathbb{C}$.
@@
If $\Gamma[s_{1}]$ already belongs to an equivalence class $\mathbb{C}\in\mathcal{C}_{EC}$, but $\Gamma[s_{2}]$ is an orphan variable—belonging to no class in $\mathcal{C}_{EC}$—then $\Gamma[s_{2}]$ is merged into $\mathbb{C}$. Swap $s_{1}$ and $s_{2}$ for the other case. If $\Gamma[s_{1}]$ and $\Gamma[s_{2}]$ belong to separate classes $\mathbb{C}_{1}$ and $\mathbb{C}_{2}$, with $\mathbb{C}_{1}\neq\mathbb{C}_{2}$, then $\mathbb{C}_{1}$ and $\mathbb{C}_{2}$ are merged into one via $\mathbb{C}\leftarrow\mathbb{C}_{1}\cup\mathbb{C}_{2}$, and one of $\mathbb{C}_{1}.\text{{rep}}$ and $\mathbb{C}_{2}.\text{{rep}}$ is deterministically elected as the representative .$\mathbb{C}.\text{{rep}}$.

### v3

@@ Abstract
Active automata learning from membership and equivalence queries is a foundational problem with numerous applications. We propose a novel variant of the active automata learning problem: actively learn finite automata using preference queries—i.e., queries about the relative position of two sequences in a total preorder—instead of membership queries. Our solution is Remap, a novel algorithm which leverages a symbolic observation table along with unification and constraint solving to navigate a space of symbolic hypotheses (each representing a set of automata), and uses satisfiability-solving to construct a concrete automaton (specifically a Moore machine) from a symbolic hypothesis. Remap is guaranteed to correctly infer the minimal automaton with polynomial query complexity under exact equivalence queries, and achieves PAC–identification ($\varepsilon$-approximate, with high probability) of the minimal automaton using sampling-based equivalence queries. Our empirical evaluations of Remap on the task of learning reward machines for two reinforcement learning domains indicate Remap scales to large automata and is effective at learning correct automata from consistent teachers, under both exact and sampling-based equivalence queries.
1 Introduction
@@
Active automata learning has applications from software engineering [37, 1] and verification [29] to interpretable machine learning [44] and learning reward machines [39, 20, 46, 16]. The classical problem formulation involves a teacher with access to a regular language and a learner which asks membership and equivalence queries to infer a finite automaton describing the regular language [5, 22].
Consider an alternative formulation: learning a finite automaton from preference and equivalence queries. [^1] Shah et al. [38] investigates choosing between membership and preference queries. A preference query resolves the relative position of two sequences in a total preorder available to the teacher. The motivation for learning from preferences stems from leveraging human preferences as a rich source of information. In fact, comparative feedback, such as preferences, is a modality which humans are apt at providing in comparison to giving specific numerical values, as shown by MacGlashan et al. [30] and Christiano et al. [12], likely due to choice overload [26]. But preferences need not be directly obtained from individual humans: preferences can also be obtained from automated systems in a frequency-based sense. For example, consider an obstacle avoidance scenario of a vehicle avoiding large debris in the middle of a roadway. Consider a dataset of a population of driver responses (vehicle trajectories) to such a scenario, procured from dashcam or traffic camera footage. The population preference of a given driving response can be automatically determined by ranking each response by the frequency of occurrence in the dataset. In fact, learning from human preference data has applications in fine-tuning language models [34], learning conditional preference networks [28, 23], inferring reinforcement learning policies [13], and learning Markovian reward functions [11, 36, 10, 27]. Possible applications for learning finite automata from preferences over sequences include inferring sequence classifications (e.g. program executions [21], vehicle maneuvers, human-robot interactions) using ordered classes (e.g. safe, risky, dangerous, fatal) with each automaton state labeled by a class, distilling interpretable preference models [44], and inferring reward machines. However, no method currently exists for learning finite automata from preferences with the termination and correctness guarantees enjoyed by classical automata learning algorithms such as $\textrm{L}^{*}$ [5].
@@
Remap [^2] Code available at https://eric-hsiung.github.io/remap/, as well as supplementary material. (Figure 1) features termination and correctness guarantees for exact and probably approximately correct [43] (PAC) identification [6] of the desired automaton, with a strong learner capable of symbolic reasoning and constraint-solving to offset the weaker preference-based signals from the teacher. While approximate learners have been proposed to learn from a combination of membership and preference queries [38], Remap is the first exact learner with formal guarantees of minimalism, correctness, and query complexity. By using unification [^3] Symbolically obtaining sets of equivalent variables from sets of equations, and substituting variables in the equations with their representatives. See Section 4.2. to navigate the symbolic space of hypotheses, and constraint-solving to construct a concrete automaton from a symbolic hypothesis, Remap identifies, in a polynomial number of queries, the minimal Moore machine isomorphic to one describing the teacher’s total preorder over input sequences when using exact equivalence queries. However, exact equivalence queries may be infeasible if the learner and teacher lack a common representation. [^4] Either a pair of identical representations, or between which a translation exists. This motivates sampling-based equivalence queries under Remap, which achieves PAC–identification. [^5] See Definition 10 for PAC–identification and related Theorems 5.4 and 5.5. Our empirical evaluations apply Remap to learn reward machines for sequential-decision making domains in the reward machine literature. We measure query complexity for the exact and PAC–identification settings, and measure empirical correctness for the PAC–identification setting.
@@
We contribute (a) Remap, a novel $\textrm{L}^{*}$ style algorithm for learning Moore machines using preference and equivalence queries; under exact and PAC-identifica-tion settings we provide (b) theoretical analysis of query complexity, correctness, and minimalism, and (c) supporting empirical results, demonstrating the efficacy of the algorithm’s ability to learn reward machines from preference and equivalence queries.
2 Background
@@
Prior to introducing Remap, we provide preliminaries on orders and finite automata, followed by a discussion of Angluin’s $\textrm{L}^{*}$ algorithm in order to highlight how Remap differs from $\textrm{L}^{*}$.
Definition 1 (Total Preorder)
[…]
Definition 2 (Total Order)
[…]
@@
Finite Automata  Automata describe sets of sequences. An alphabet $\Sigma$ is a set whose elements can be used to construct sequences; $\Sigma^{*}$ represents the set of sequences of any length created from elements of $\Sigma$. A sequence $s\in\Sigma^{*}$ has integer length $|s|\geq 0$; if $|s|=0$, then $s$ is the empty sequence $\varepsilon$. An element $\sigma\in\Sigma$ has length 1; if $s,t\in\Sigma^{*}$, then $s\cdot t$ represents $s$ concatenated with $t$, with length $|s|+|t|$. Different types of finite automata have different semantics. Deterministic automata feature a deterministic transition function $\delta$ defined over a set of states $Q$ and an input alphabet $\Sigma^{I}$. An output alphabet $\Sigma^{O}$ may be present to label states or transitions using a labeling function $L$. The tuple $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ is a Moore machine [33], where $q_{0}\in Q$ is the initial state, $\delta:Q\times\Sigma^{I}\rightarrow Q$ describes transitions, and $L:Q\rightarrow\Sigma^{O}$ associates outputs with states. Extended, $\delta:Q\times(\Sigma^{I})^{*}\rightarrow Q$, where $\delta(q,\varepsilon)=q$ and $\delta(q,\sigma\cdot s)=\delta(\delta(q,\sigma),s)$. In Mealy machines [32], outputs are associated with transitions instead, so $L:Q\times\Sigma^{I}\rightarrow\Sigma^{O}$. Mealy and Moore machines are equivalent [19] and can be converted between one another. Reward machines are an application of Mealy machines in reinforcement learning and are used to express a class of non-Markovian reward functions.
@@
In Angluin’s seminal active learning algorithm, $\textrm{L}^{*}$ [5], the learner learns $\hat{f}$ as a binary classifier, where $|\Sigma^{O}|=2$, to determine sequence membership of a regular language by querying $\mathcal{T}$ with: (i) membership queries, where the learner asks $\mathcal{T}$ for the value of $f(s)$ for a particular sequence $s$, and (ii) equivalence queries, where the learner asks $\mathcal{T}$ to evaluate whether for all $s\in(\Sigma^{I})^{*}$, $\hat{f}(s)=f(s)$. For the latter query, $\mathcal{T}$ returns True if the statement holds; otherwise a counterexample $c$ for which $\hat{f}(c)\neq f(c)$ is returned. An observation table $\langle S,E,T\rangle$ records the concrete observations acquired by the learner’s queries (see Figure 2). Here, $S$ is a set of prefixes, $E$ is a set of suffixes, and $T$ is the empirical observation function that maps sequences to output values—$T:(S\cup(S\cdot\Sigma^{I}))\cdot E\rightarrow\Sigma^{O}$. The observation table is a two-dimensional array; $s\in(S\cup(S\cdot\Sigma^{I}))$ indexes rows, and $e\in E$ indexes columns, with entries given by $T(s\cdot e)$. Any proposed hypothesis $\hat{f}$ must be consistent with $T$. The algorithm operates by construction: a deterministic transition function must be found exhibiting consistency (deterministic transitions) and operates in a closed manner over the set of states. If the consistency or closure requirements are violated, then membership queries are executed to expand the observation table. Once a suitable transition function is found, a hypothesis $\hat{f}$ can be made and checked via the equivalence query. The algorithm terminates if $\hat{f}(s)=f(s)$ for all $s\in(\Sigma^{I})^{*}$; otherwise $\textrm{L}^{*}$ continues on by adding counterexample $c$ and all its prefixes to the table, and finds another transition function satisfying the consistency and closure requirements.
Figure 2: $\textrm{L}^{*}$ example. Figure 3 illustrates Remap. $\textrm{L}^{*}$ records concrete values from membership queries in observation table. Green symbolizes new information. Colors in (d) visually highlight transition inconsistencies.
@@
Consequently, we consider how $\textrm{L}^{*}$-style learning can be used to learn Moore machines from preference queries over sequences, as a foray into understanding how finite automaton structure can be learned from comparison information.
3 Problem Statement
@@
We consider how a Moore machine $\langle Q,q_{0},\Sigma^{I},\Sigma^{O},\delta,L\rangle$ can be actively learned from preference queries over $(\Sigma^{I})^{*}$. We focus on the case of a finite sized $\Sigma^{I}$ and $\Sigma^{O}$ known to both teacher and learner. With a preference query, the learner asks the teacher $\mathcal{T}$ which of two sequences $s_{1}$ and $s_{2}$ is preferred, or if both are equally preferable. This requires a preference model, which we assume represents a total preordering over $(\Sigma^{I})^{*}$; we also assume $\Sigma^{O}$ is totally ordered. Thus, we consider a preference model for $\mathcal{T}$ where $f:(\Sigma^{I})^{*}\rightarrow\Sigma^{O}$ is consistent with both orderings, i.e., $\mathcal{T}$ prefers $s_{1}$ over $s_{2}$ if $f(s_{1})>f(s_{2})$, or otherwise has equal preference if $f(s_{1})=f(s_{2})$.
Several options for evaluating hypothesis equivalence (is $\hat{f}\equiv f$?) can be defined for this problem formulation. We first review the definition of exact equivalence, followed by an alternative notion of equivalence which respects ordering:
Definition 3 (Exact Equivalence)
Given a hypothesis $\hat{f}$ and a reference $f$, $\hat{f}$ is exactly equivalent to $f$ if $\hat{f}(s)=f(s)$ for all $s$ in $(\Sigma^{I})^{*}$.
Definition 4 (Order-Respecting Equivalence)
Given a hypothesis $\hat{f}$ and a reference $f$, $\hat{f}$ is order-respecting equivalent to $f$ if, for all $s,t$ in $(\Sigma^{I})^{*}$, there exists a relation $\mathbf{R}_{s,t}\in\{=,>,<\}$ such that $\hat{f}(s)\mathbf{R}_{s,t}\hat{f}(t)\iff f(s)\mathbf{R}_{s,t}f(t)$.
Since a hypothesis $\hat{f}$ satisfying exact equivalence is also always order-respect-ing, we focus on exact equivalence queries where $\mathcal{T}$ returns feedback along with counterexample $c$.
Definition 5 (Exact Equivalence Query with Feedback)
Given a hypothesis $\hat{f}$ and a reference $f$, an exact equivalence query EQ returns the following triple:
$\text{{\emph{EQ}}}(\hat{f})=\langle\forall s\in(\Sigma^{I})^{*}:\hat{f}(s)=f(s),c:\exists c\text{ \emph{s.t.} }\hat{f}(c)\neq f(c),\phi(c)\rangle$
where $c$ is a counterexample if one exists, and $\phi(c)$ is feedback associated with the counterexample, interpreted as a constraint.
@@
The strength of feedback $\phi$ directly impacts how many hypotheses per counterexample can be eliminated by the learner: returning $\phi(c):=\hat{f}(c)\neq f(c)$, which means the value $f(c)$ is not $\hat{f}(c)$, is weak feedback compared to returning $\phi(c):=\hat{f}(c)=f(c)$, which means the value of $\hat{f}(c)$ should be $f(c)$, or returning $\phi(c):=\hat{f}(c)\in X$ (where $X\subset\Sigma^{O}$, which means the value of $\hat{f}(c)$ should be in a subset of $\Sigma^{O}$). Although Remap outputs a concrete Moore machine, Remap navigates symbolic Moore machine666i.e., a Moore machine with symbolic values as outputs. space using solely preference information, while concrete information assists in selecting the concrete hypothesis. Thus, our theoretical analysis considers equivalence queries that provide counterexamples with strong feedback: $\phi(c):=\hat{f}(c)=f(c)$.
4 The Remap Algorithm
@@
Figure 3: Remap example. Figure 2 illustrates $\textrm{L}^{*}$. Remap performs preference queries and records variables in symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$ within a SymbolicFill. (a) initializes $\langle S,E,T;\mathcal{C},\Gamma\rangle$; (b) expands $S$ with sequence $b$, followed by SymbolicFill; yields unified, closed, and consistent $\langle S,E,T;\mathcal{C},\Gamma\rangle$; (c) MakeHypothesis yields concrete hypothesis $h_{1}$ from symbolic hypothesis and constraints in $\mathcal{C}$; (d) submits $h_{1}$ via equivalence query; receives and processes counterexample $bab$ with feedback $f(bab)=0$, adding $bab$, $ba$, and $b$ to $S$, performs a SymbolicFill, sets the value of equivalence class for $T(bab)$ to $f(bab)$; yields an inconsistent table. (e) shows resulting consistent table. Figure 2f shows ground truth Moore machine with $\Sigma^{I}=\{a,b\},\Sigma^{O}=\{0,1\}$. Green symbolizes new information. Colors in (d) visually highlight transition inconsistencies.
@@
Central to Remap are four core components: (1) a new construct called a symbolic observation table (Section 4.1), shown in Figure 3a, 3b, 3d, and 3e; (2) as well as an associated algorithm for unification [35] (Section 4.2) inspired by Martelli and Montanari [31] to contain the fresh variable explosion; both enable the learner to (3) generate symbolic hypotheses (Section 4.3) purely from observed symbolic constraints, along with (4) a constraint solver for obtaining a concrete hypothesis (Section 4.4). We first discuss the four components and how they fit together into Remap (Section 4.5). Afterwards, we illustrate the correctness and termination guarantees.
4.0.1 4.1 Symbolic Observation Tables
@@
Recall that in classic $\textrm{L}^{*}$, the observation table entries are concrete values obtained from membership queries (Figure 2a). However, observations obtained from preference queries are constraints, rather than concrete values, indicating that for a pair of sequences $s_{1}$ and $s_{2}$, one of $f(s_{1})>f(s_{2})$, $f(s_{1})<f(s_{2})$, or $f(s_{1})=f(s_{2})$ holds. We therefore introduce the symbolic observation table $\langle S,E,T;\mathcal{C},\Gamma\rangle$, where $S$ is a set of prefixes, $E$ is a set of suffixes, $\mathcal{C}$ is the set of known constraints, $\Gamma$ is a context which uniquely maps sequences to variables, with the set of variables $\mathcal{V}$ in the context $\Gamma$ given by $\mathcal{V}=\{\Gamma[s\cdot e]|s\cdot e\in(S\cup(S\cdot\Sigma^{I}))\cdot E\}$, and $T:(S\cup(S\cdot\Sigma^{I}))\cdot E\rightarrow\mathcal{V}$ maps queried sequences to variables. Thus, in a symbolic observation table, the entry for each prefix-suffix pair is a variable, rather than a concrete value; and the constraints over those variables are stored in $\mathcal{C}$ (Figure 3a, 3b, 3d, 3e). The constraints from the preferences of $\mathcal{T}$ over $f$ about $s_{1}$ and $s_{2}$ correspond with $T(s_{1})>T(s_{2})$, $T(s_{1})<T(s_{2})$, or $T(s_{1})=T(s_{2})$, respectively.
Definition 6
@@
An equivalence class $\mathbb{C}$ of variables is a set with the property that all members of $\mathbb{C}$ are equivalent to each other. The representative $\llbracket\mathbb{C}\rrbracket$ of $\mathbb{C}$ is a deterministically elected member of $\mathbb{C}$. The set of variables $\mathcal{V}$ can be partitioned into disjoint equivalence classes. The set of equivalence classes $\mathcal{C}_{EC}$ corresponds with the partitioning of $\mathcal{V}$ into the smallest possible number of equivalence classes consistent with equality constraints. Let =$\mathcal{R}=\{\llbracket\mathbb{C}\rrbracket|\mathbb{C}\in\mathcal{C}_{EC}\}$ be the set of representatives.
4.0.2 4.2 Unification and Constraints
@@
Since preference queries return observations comparing the values of two sequences, we leverage a simple unification algorithm to ensure the number of unique variables in the table remains bounded by $|\Sigma^{O}|$. As a reminder, unification involves symbolically solving for representatives and rewriting all constraints and variables in terms of those representatives. Whenever we observe the constraint $f(s_{1})=f(s_{2})$, this implies $T(s_{1})=T(s_{2})$, so we add $\Gamma[s_{1}]$ and $\Gamma[s_{2}]$ to an equivalence class $\mathbb{C}$, and elect a representative from $\mathbb{C}$. The unification algorithm is presented in Appendix Algorithm 9.
@@
If $\Gamma[s_{1}]$ already belongs to an equivalence class $\mathbb{C}\in\mathcal{C}_{EC}$, but $\Gamma[s_{2}]$ is an orphan variable—belonging to no class in $\mathcal{C}_{EC}$—then $\Gamma[s_{2}]$ is merged into $\mathbb{C}$. Swap $s_{1}$ and $s_{2}$ for the other case. If $\Gamma[s_{1}]$ and $\Gamma[s_{2}]$ belong to separate classes $\mathbb{C}_{1}$ and $\mathbb{C}_{2}$, with $\mathbb{C}_{1}\neq\mathbb{C}_{2}$, then $\mathbb{C}_{1}$ and $\mathbb{C}_{2}$ are merged into one via $\mathbb{C}\leftarrow\mathbb{C}_{1}\cup\mathbb{C}_{2}$, and one of $\llbracket\mathbb{C}_{1}\rrbracket$ and $\llbracket\mathbb{C}_{2}\rrbracket$ is deterministically elected as the representative $\llbracket\mathbb{C}\rrbracket$.